OpenStudy (thedj4jc):

Which of the following equations is quadratic? x(x^2 + 1) = 0 5(4x + 2) = 3 (x + 3)(x + 4) = 5

OpenStudy (akashdeepdeb):

Expand each of them and whichever has a degree of 2 (quadratic) will be your answer.

OpenStudy (thedj4jc):

What do you mean?

OpenStudy (thedj4jc):

@AkashdeepDeb

OpenStudy (thedj4jc):

@SolomonZelman

OpenStudy (astrophysics):

degree two meaning ax^2...etc

OpenStudy (astrophysics):

\[x^2 \implies degree ~ 2\] you can think of it as degree two since it has 2 on the x I guess.

OpenStudy (thedj4jc):

What do you mean "degree 2?"

OpenStudy (astrophysics):

I just explained it

OpenStudy (astrophysics):

So expand your problem and if it has x^2 that's the one..

OpenStudy (thedj4jc):

So, A????

OpenStudy (astrophysics):

Just expand it.

OpenStudy (astrophysics):

Each of the problems, meaning multiply through if that helps.

OpenStudy (astrophysics):

What is \[x \times x^2\]

OpenStudy (thedj4jc):

x^3, I think.

OpenStudy (astrophysics):

Yeah, so x(x^2+1)= 0 means x^3+x=0

OpenStudy (astrophysics):

Do that with the others.

OpenStudy (astrophysics):

And notice it's x^3 so it's not of degree 2.

OpenStudy (thedj4jc):

20x + 10 = 3 is B

OpenStudy (astrophysics):

Good

OpenStudy (astrophysics):

Last one?

OpenStudy (astrophysics):

You can foil that out, meaning